# Holistic Optimization Framework for FPGA Accelerators

### v2

@@ IV NLP Formulation
Table I delineates the sets, variables, and constants utilized in our NLP formulation.
Constants Description
@@
$p_{i}^{l}$ Position of the loop $l$ under the $i$-th permutation
@@ 8 F[i1][j1] += C[i][k] * D[k][j1];}
Bit Width  $\mathcal{B}$ denotes the number of elements that can be transferred simultaneously, determined by the bit width and the data type. Hence, if we have a bit width under 512 bits for float the set is $\{1,2,4,8,16\}$. Equation 3 computes the bit width for each array based on the last dimension of the data-tile transferred on-chip. For example, the array D in Fused task 1 is transferred in line 15 with a size of $222\times 32$, so $S^{last}_{D}=32$, which is divisible by 16. Therefore, it has a bit width of 16.
@@
Permutation  Equation 4 requires the NLP to choose identical permutations for loops that are shared by statements fused within the same task. For example, in Fused task 1, the loops iterating statement S2 can be permuted as ($i0$, $j0$) or ($j0$, $i0$), and similarly, loops iterating statement S3 can be permuted as ($i0$, $j0$) or ($j0$, $i0$). However, since the loops $i0$ and $j0$ iterate both S2 and S3, they must use the same permutation; either ($i0$, $j0$) for both or ($j0$, $i0$) for both.
@@
DSP Utilization  Equation 10 constrains the number of DSPs used based on the available DSP resources. In contrast to [58, 57, 59], we utilize pessimistic DSP utilization. This approach is necessary because concurrent execution and resource reuse between tasks are not feasible when two tasks can run simultaneously. Given $DSP_{+}=2$, $DSP_{*}=3$, and $II_{S3}=3$, the DSP usage for Task 3 is calculated as $(2+3)\times 1824$, accounting for the unroll factor. However, since the loop is pipelined with $II=3$, the HLS compiler optimizes resource usage, effectively reducing the DSP count by approximately dividing by $II$.
@@
SLR Selection  For each task, the NLP determines the SLR on which the task will be implemented (Equation 11). Additionally, Equations 7 and 10 are applied to each SLR to manage resource allocation per SLR effectively.
@@ V Code Generation
Intra-Task  Each intra-task, corresponding to an intra-tile selected by the NLP, is implemented as a fully unrolled, independent function without communication with off-chip memory. Data for these tasks resides entirely on-chip.
If the task involves reduction loops, inter-tile reduction loops are integrated into the function and pipelined. While the pipeline initiation interval (II) is greater than 1 due to reduction dependencies, other operations in the statement are pipelined efficiently.
Padding is handled at the intra-tile level. Non-reduction loops remain unchanged, allowing computation of padding values without excessive resource usage. For reduction loops, full tiles are computed first, and the intra-tile loop is adjusted to handle padding for partial tiles accurately.
Inter-Tile Loop  For each fused task and each inter-tile loop, we generate independent functions to facilitate efficient double or triple buffering. Using information from the NLP, we determine which arrays are defined and which are transferred at each level of granularity. Once we identify the data being transferred, we implement double buffering if only reads or writes are involved, or triple buffering if both reads and writes are required. This strategy allows us to overlap the operations of reading, storing, and computing at the innermost level, optimizing both communication and computation overlap.
Concurrent Execution  Using the dataflow pragma, independent tasks execute concurrently without requiring manual rewrites. Computation in a receiving task begins as soon as its shared buffer contains sufficient data.
SLR Management  The NLP determines the SLR ID for each task, specifying where it will be executed. Prometheus generates a separate C++ file for each SLR, and data transfers between SLRs are managed via ap_axiu streams, ensuring efficient communication and minimizing transfer overhead.
@@ VI-A Setup and Experimental Evaluation
The objective of this evaluation is to demonstrate the capability of our framework to generate code with high QoR, whether for memory-bound or computation-bound kernels. To achieve this, we compare our work with Allo, ScaleHLS, Sisyphus, and AutoDSE. We do not include a comparison with Stream-HLS [9] due to the unavailability of its source code. For ScaleHLS [78] and Allo [15], we apply the same method as our framework by padding memory transfers to attain the maximal bit width (512 bits) from off-chip to on-chip memory. Their kernels assume data are already in on-chip memory, so we modified their code to include off-chip to on-chip data transfers. Allo’s 2mm and 3mm kernels already handle this transfer, requiring no modifications. We left Sisyphus and AutoDSE unchanged, as they already optimize bit width according to the problem size. We use AutoDSE with the bottleneck method, setting a DSE timeout of 1,000 minutes and an HLS synthesis timeout of 180 minutes per task. For Allo-generated designs, we utilize kernels from their artifact repository since Allo does not employ a DSE [1]. Sisyphus generated designs are generated using parameters consistent with the paper [59], and ScaleHLS designs are compiled using their open-source compiler [2]. For RTL simulation, we assume that the frameworks can utilize all the resources on the U55C FPGA with a constraint of partitioning any array of 1,024 due to AMD/Xilinx limitations. For on-board FPGA evaluations, we consider two scenarios. The first scenario utilizes 60% of one Super Logic Region (SLR), equivalent to 20% of total board resources, for all frameworks. The second scenario is unique to our framework, leveraging all three SLRs, with 60% utilization per SLR. Most frameworks are not place-and-route aware, leading to congestion issues that prevent bitstream generation. Limiting frameworks to one SLR increases the likelihood of meeting timing requirements. AutoDSE, for instance, lacks dataflow support and applies pragmas within a single function, making multi-SLR bitstream generation impossible without manual intervention to split the function across SLRs. If congestion persists within an SLR, we adjust the NLP constraints to address the issue and regenerate the HLS-C++ file in just a few minutes.
@@ VI-B Comparison
The average time to solve the NLP and generate the designs presented in this section is 31.70 seconds, with a maximum of 178.41 seconds for 3mm with 3 SLRs.
@@
TABLE II: Comparison of the throughput (GF/s) with RTL Simulation
@@
Table IV shows the parameters found by the NLP. We use the same name of the iterator as Polybench 4.2.1 [55] code. $Si$ represent the statement in position $i$ in the code. The second column gives the statement fused inside the same tasks, the third column sets the fused task order and loop order found by the NLP for the fused task and the last column supplies the data-tile size found by the NLP, if the array is present in a different fused task the fused-task (defined in the second column) is precise.
@@ VII Related Work
Dataflow  Dataflow principles have been extensively studied in models such as Kahn Process Networks [34], Dennis Dataflow [23], synchronous dataflow languages [41, 12], and for FPGA applications [52, 13, 3, 26, 11, 5, 74, 15]. DaCe [11] introduces Stateful DataFlow multiGraphs to separate program definition from optimization, enabling transformations like tiling and double-buffering, though requiring user intervention. Stream-HLS [9] automatically generates dataflow; however, it assumes that data are already on-chip, thereby overlooking communication with off-chip memory. Additionally, it offers very limited opportunities for parallelism. Flower [5] automates FPGA dataflow development but limits parallelism. Frameworks like [15, 74], built on HeteroCL [37], optimize data placement and compute scheduling for heterogeneous systems, maximizing data reuse and bandwidth utilization. Systolic arrays [70, 21, 33, 38, 10] offer efficient computation for specific patterns but lack generalization. Application-specific frameworks [17, 29, 50, 22, 62, 24, 35, 6] demonstrate dataflow advantages but do not generalize across domains. RapidStream-TAPA [27, 28] enhances the performance of dataflow designs and automates SLR placement. However, it requires an optimized kernel with a dataflow structure as input.
@@ VIII Conclusion
In this work, we introduce Prometheus, a comprehensive FPGA performance enhancement framework. By integrating techniques like task fusion, automatic tiling, loop permutation, and overlap of communication and computation, Prometheus is intended to minimize latency within stringent resource constraints. Our methodology utilizes NLP to navigate complex design spaces and automate parameter discovery, achieving superior QoR compared to existing tools.

### v3

@@ IV NLP Formulation
Constant Description
@@
$p_{i}^{l}$ Position of the permutation $i$ for the loop $l$
@@ 8 F[i1][j1] += C[i][k] * D[k][j1];}
Bit Width  $\mathcal{B}$ represents the possible bit width for the data type. Hence, if we have a bit width under 512 bits for float the set is $\{1,2,4,8,16\}$. Equation 3 computes the bit width for each array based on the last dimension of the data-tile transferred on-chip. For example, the array D in Fused task 1 is transferred in line 15 with a size of $222\times 32$, so $S^{last}_{D}=32$, which is divisible by 16. Therefore, it has a bit width of 16.
@@
Permutation  Equation 4 ensures that the NLP selects a unique permutation for each loop that iterates through every statement. Equation 5 requires the NLP to choose identical permutations for loops that are shared by statements fused within the same task. For example, in Fused task 1, the loops iterating statement S2 can be permuted as ($i0$, $j0$) or ($j0$, $i0$), and similarly, loops iterating statement S3 can be permuted as ($i0$, $j0$) or ($j0$, $i0$). However, since the loops $i0$ and $j0$ iterate both S2 and S3, they must use the same permutation; either ($i0$, $j0$) for both or ($j0$, $i0$) for both.
$\forall s\in\mathcal{S},\forall i\in\mathcal{P}_{s},\sum_{p}p_{i}^{l}=1$ (4)
@@
DSP Utilization  Equation 11 constrains the number of DSPs used based on the available DSP resources. In contrast to [59, 58, 60], we utilize pessimistic DSP utilization. This approach is necessary because concurrent execution and resource reuse between tasks are not feasible when two tasks can run simultaneously.
Given $DSP_{+}=2$, $DSP_{*}=3$, and $II_{S3}=3$, the DSP usage for Task 3 is calculated as $(2+3)\times 1824$, accounting for the unroll factor. However, since the loop is pipelined with $II=3$, the HLS compiler optimizes resource usage, effectively reducing the DSP count by approximately dividing by $II$.
@@
SLR Selection  For each task, the NLP determines the SLR on which the task will be implemented (Equation 12). Additionally, Equations 8 and 11 are applied to each SLR to manage resource allocation per SLR effectively.
@@ V Code Generation
Intra-Task  Each intra-task, corresponding to the intra-tile selected by the NLP, is implemented as an independent, fully unrolled function with no communication with off-chip memory. Data for these tasks resides on-chip. If the task includes reduction loops, inter-tile reduction loops are incorporated and pipelined within the function. The pipeline’s II is strictly greater than 1 due to reduction loop dependencies, but other operations within the statements are effectively pipelined. Padding is managed at the intra-tile level. For non-reduction loops, padding does not alter the intra-task, allowing computation of padding values without unnecessary resource usage for minor latency improvements. In contrast, for reduction loops, full tile computations are performed first, followed by adjusting the intra-tile loop to handle padding and compute partial tiles accurately.
Inter-Tile Loop  For each fused task and each inter-tile loop, we generate independent functions to facilitate efficient double or triple buffering. Using information from the NLP, we determine which arrays are defined and which are transferred at each level of granularity. Once we identify the data being transferred, we implement double buffering if only reads are involved, or triple buffering if both reads and writes are required. This strategy allows us to overlap the operations of reading, storing, and computing at the innermost level, optimizing both communication and computation overlap.
Concurrent Execution  With pragma dataflow, independent tasks run concurrently without rewriting. Computation begins as soon as the receiving task’s shared buffer has enough data.
SLR  The NLP determines the SLR on which each task will be executed. A separate C++ file is generated for each SLR, and data transfers between SLRs are handled using ap_axiu.
@@ VI-A Setup and Experimental Evaluation
The objective of this evaluation is to demonstrate the capability of our framework to generate code with high QoR, whether for memory-bound or computation-bound kernels. To achieve this, we comare our work with Allo, ScaleHLS, Sisyphus, and AutoDSE. We execute the designs for these frameworks using identical parameters as outlined in Section VI-A. For ScaleHLS [79] and Allo [14], we apply the same method as our framework by padding memory transfers to attain the maximal bit width (512 bits) from off-chip to on-chip memory. Their kernels assume data are already in on-chip memory, so we modified their code to include off-chip to on-chip data transfers. Allo’s 2mm and 3mm kernels already handle this transfer, requiring no modifications. We left Sisyphus and AutoDSE unchanged, as they already optimize bit width according to the problem size. We use AutoDSE with the bottleneck method, setting a DSE timeout of 1,000 minutes and an HLS synthesis timeout of 180 minutes per task. For Allo-generated designs, we utilize kernels from their artifact repository since Allo does not employ a DSE [1]. Sisyphus generated designs are generated using parameters consistent with the paper [60], and ScaleHLS designs are compiled using their open-source compiler [2]. For RTL simulation, we assume that the frameworks can utilize all the resources on the U55C FPGA with a constraint of partitioning any array of 1,024 due to AMD/Xilinx limitations. For on-board FPGA evaluations, we consider two scenarios. The first scenario utilizes 60% of one Super Logic Region (SLR), equivalent to 20% of total board resources, for all frameworks. The second scenario is unique to our framework, leveraging all three SLRs, with 60% utilization per SLR. Most frameworks are not place-and-route aware, leading to congestion issues that prevent bitstream generation. Limiting frameworks to one SLR increases the likelihood of meeting timing requirements. AutoDSE, for instance, lacks dataflow support and applies pragmas within a single function, making multi-SLR bitstream generation impossible without manual intervention to split the function across SLRs. If congestion persists within an SLR, we adjust the NLP constraints to address the issue and regenerate the HLS-C++ file in just a few minutes.
@@ VI-B Comparison
The average time to solve the NLP and generate the designs presented in this section is 31.7 seconds, with a maximum of 178.416 seconds for 3mm with 3 SLRs.
@@
TABLE II: Comparison of the throughput (GF/s) of various optimization framework with RTL Simulation
@@
Table IV shows the parameters found by the NLP. We use the same name of the iterator as Polybench 4.2.1 [56] code. $Si$ represent the statement in position $i$ in the code. The second column gives the statement fused inside the same tasks, the third column sets the fused task order and loop order found by the NLP for the fused task and the last column supplies the data-tile size found by the NLP, if the array is present in a different fused task the fused-task is precise.
@@ VII Related Work
Dataflow  Dataflow principles have been extensively studied in models such as Kahn Process Networks [34], Dennis Dataflow [22], synchronous dataflow languages [42, 11], and for FPGA applications [53, 12, 3, 25, 10, 5, 75, 14]. DaCe [10] introduces Stateful DataFlow multiGraphs to separate program definition from optimization, enabling transformations like tiling and double-buffering, though requiring user intervention. Flower [5] automates FPGA dataflow development but limits parallelism. Frameworks like [14, 75], built on HeteroCL [38], optimize data placement and compute scheduling for heterogeneous systems, maximizing data reuse and bandwidth utilization. Systolic arrays [71, 20, 33, 39, 9] offer efficient computation for specific patterns but lack generalization. Application-specific frameworks [16, 29, 51, 21, 63, 23, 36, 6] demonstrate dataflow advantages but do not generalize across domains.
@@ VIII Conclusion
In this work, we introduce Prometheus, a comprehensive FPGA performance enhancement framework. By integrating techniques like task fusion, automatic tiling, loop permutation, and overlap of communication and computation, Prometheus is intended to maximize throughput within stringent resource constraints. Our methodology utilizes NLP to navigate complex design spaces and automate parameter discovery, achieving superior QoR compared to existing tools.